6x+2y-3z=-17;7x-5y+z=72;2x+8y+3z=-21
solve by elimination

Answer by Alan3354(69443)

6x+2y-3z=-17
7x-5y+z=72
2x+8y+3z=-21
solve by elimination
=============
Multiply the 2nd eqn by 3, then eliminate z.
You'll have 2 eqns in x & y.